This will use distance velocity and time and will require 2 equations
I bicycle at 20 mph and leave the starting point at 8 AM
My buddy drives at 40 mph and leaves the starting point at 10 AM following me
How far have I gone when he catches up and how much time has passed
I got the answer as, I will have traveled 80 miles and 4 hours will have passed, but I would like to see it in a formula, so I can understand the algebra formula

Answer by scott8148(6628)

if your time is T, then your buddy's time is T-2 (two hours less)
you know the rates
when he catches up, you both will have travelled an EQUAL distance
d = r * t

20 * T = 40 * (T - 2) ___ 20T = 40T - 80 ___ -20T = -80 ___ T = 4

substituting back to find distance ___ d = 20 * 4 ___ d = 80